OpenStudy (anonymous):

What is the equation of a line that has a slope of -5 and passes through the point (0, 2)? a) y=-5x b) y=-5x+2 c) y=2x-5

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\]

pooja195 (pooja195):

We need to use this formula it is called the point slope formula m=slope (x1,y1)

OpenStudy (anonymous):

okay so, so y=mx+b right... so in this case y=-5x + b

OpenStudy (anonymous):

plug in values for x and y of your point in this equation

OpenStudy (anonymous):

(2=-5(0)+b)

OpenStudy (anonymous):

solve for b

OpenStudy (anonymous):

b=2

OpenStudy (anonymous):

so y=-5x +2

pooja195 (pooja195):

Or just give the answer -_-